Net Sales 46021, 29865, 37536, 36012, 34553 (2017-2013)

  1. Determine the sales-growth rate (in percentage) from 2014 to 2017. What trend do we observe?
  2. If we were to use the four-year average sales growth to forecast 2018 sales, what rate would we use?
  3. c. The annual report reveals that 2016 included a 53rd week (as is common for retailers, every four years) and that the company acquired a smaller company during that year. Sales for the newly acquired company were included in the consolidated income statement: from the date of acquisition, which amounted to $5,235 million. Use this information to refine the sales-growth rate. (Hint: Consider multiplying sales for a 53-week year by 52/53 to get an apples-to-apples comparison with numbers from 52-week years.)
Answer #1

Net Sales for the year 2013 = $34,553

2014 = $36,012

2015 = $37,536

      2016 = $29,865

   2017 = $46,021

Assume that all the above figures given are in millions

(a) Sales Growth Rate

Sales Growth Rate = (Current year Sales - Previous year Sales) / Previous year Sales * 100

Growth rate In the year 2014 = ($36,012 - $34,553) / $34,553

= $1,459 / $34,553 * 100

   = 0.0422 * 100

   = 4.22%

Growth rate In the year 2015 = ($37,536 - $36,012) / $36,012 * 100

= $1,524 / $36,012 * 100

= 0.0423 * 100

  = 4.23%

Growth rate In the year 2016 = ($29,865 - $37,536) / $37,536 * 100

= - $7,671 / $37,536 * 100

   = - 0.2043 * 100

= - 20.43%

Growth rate In the year 2017 = ($46,021 - $29,865) / $29,865* 100

= $16,156 / $29,865 * 100

  = 0.5409 * 100

= 54.09%

We can observe that in 2014 and 2015 there is incremental improvement in the sales growth rate, but in 2016 there is decline in the growth rate which is even negative but the company has recovered from that situation as there is positive improvement in sales growth rate in the year of 2017.

(b) Four years average growth rate

   = (Growth rate in 2014+Growth rate in 2015+Growth rate in 2016+Growth rate in 2017) / Total no. of years

= (4.22%+4.23%-20.43+54.09) * 4

= 42.11 / 4

= 10.52% (Approx.)

For forecasting the sales of 2018, if four year average rate is used then 10.52% will be appropriate.

(c) Given that the sales of 2016 included sales of 53 weeks instead of 52 weeks and it also includes $5,235 million sales of a company which is recently acquired during the year.

So the $5,235 should be excluded from the given sales figures

= $29,865 - $5,235

= $24,630

In 2016, $24,630 are sales for 53 weeks then sales for 52 weeks = 52/53 * $24,630

   Revised Sales for 2016 = $24,165

Revised growth rate for 2016 = (Current year revised Sales - Previous year Sales) / Previous year Sales * 100

= ($24,165 - $37,536 / $37,536* 100

= - $13,371/ $37,536 * 100

  = - 35.62%

Revised growth rate for 2017 = (Current yr Sales - Previous yr revised Sales)/Previous year revised Sales *100

= ($46,021 - $24,165) / $24,165 * 100

= $21,856/ $24,165 * 100

  = 90.44%